The price of Mega Polls, Inc. stock is $30.20 per share. The forecast is that the stock has a 60% chance of rising to $35.25 or a 40% chance of falling to $27.10. A call option on Mega’s stock has an exercise price of $28 per share. If the Treasury bill rate is 3.0%, what is the likely price of a 1-year call option? (Round your answer to 2 decimal places.)

Call Price?

*Note all information is provided no other info available
Answer #1

The likely price of a 1 year call option is calculated as follows:

If the stock price goes to $ 35.25, then value of call option = Max (0, $ 35.25 - $ 28 )

= 7.25

If the stock price goes to $ 27.10, then value of call option = Max ( 0, $ 27.10 - $ 28 )

= 0

Now the up move probability is 60% and probability of down move is 40%, so the likely price of a 1 year call option is

= 7.25 x 60% + 0 x 40%

= $ 4.35

Now in case if we want to compute the present value of the call option we can get it by discounting it at 3% i.e.

= 4.35 / 1.03

= $ 4.22 Approximately

Feel free to ask in case of any query relating to this question